p⃗ = mv⃗

\begin{aligned}p_{x} & = mv_{x} \\ p_{y} & = mv_{y} \\ p_{z} & = mv_{z} \end{aligned}

momentum : = 1 \frac{kg m}{s}

KE : 1 Joule = 1 \frac{kg m^{2}}{s^{2}}

(momentum)(velocity) = (\frac{kg m}{s})(\frac{m}{s}) = \frac{kg m^{2}}{s^{2}} = (energy)

∑ mv⃗_{initial} = ∑ mv⃗_{final}

\begin{aligned}∑ mv_{x,i} & = ∑ mv_{x,f} \\ ∑ mv_{y,i} & = ∑ mv_{y,f} \\ ∑ mv_{z,i} & = ∑ mv_{z,f} \end{aligned}